Point charges q1=50uC and q2=—25uC are placed 1.0 m apart. 


(a) What is the electric field at a point midway between them? 

(b) What is the force on a Charge q3=20uC situated there? 



1
Expert's answer
2021-05-30T13:29:06-0400

a)


"E=\frac{72}{20\cdot10^{-6}}=3.6\cdot10^{6}\frac{V}{m}"

b)


"F_3=k\frac{q_1q_3}{r^2}+k\frac{q_2q_3}{r^2}\\
F_3=9\cdot10^{9}\frac{(50\cdot10^{-6})(20\cdot10^{-6})}{0.5^2}\\+9\cdot10^{9}\frac{(50\cdot10^{-6})(20\cdot10^{-6})}{0.5^2}=72\ N"